Signature Name& Surname: Student No 02 (40 pts). Residual stresses generated because of cooling from high curing temperatures: A [o laminate of AS-4 carbon fiber-epoxy is cured at temperature T 180 °C and slowly cooled 902 20 °C. Determine the residual stresses generated in each layer because of cooling from the Assume each layer in the laminate has a thickness to. The composite has the following curing temperature. thermomechanical properties Et G12 V12 α11 (per °C) -1.8 x10-6 α22 (GPa) (GPa) (GPa) (per °C) 142 27x 10
Answer #1

Given, betore Covting to epoxy cash n tibu la expound and called edal them et inihal uglin z 2981x106 -4320 1X Io 142xi0 riceAnd also we kvver Eil = 트4-142 (2 E22 10.3 142-13-18 14202-6740 2f40 = 6740 → 10.3(I3X)红+ ん= 23.73 MPou